OpenStudy (anonymous):

If GB = GE and AC = 12 cm, find the length of DF. answers : 15 cm 6 cm 12 cm 18 cm

OpenStudy (anonymous):

http://tinypic.com/r/2mfg2th/5

OpenStudy (anonymous):

i think its 6 ?

OpenStudy (loser66):

I think it's 12

OpenStudy (anonymous):

is it because theyre the same lines?

OpenStudy (loser66):

nope, when I apply Pythagorean theorem, I have the same answer for them let radius of the circle is r, let GB = GE =a right triangle GBA has AB^2 = r^2 -a^2 and right triangle GEF has ED^2 = r^2 -a^2, too therefore, AB = EF and B and E are middle points of AC and DF respectively therefore AC=DF =12

OpenStudy (anonymous):

ohh so you have to use the pythagorean theorem? okay okay thanks(:

OpenStudy (loser66):

yw